OpenStudy (anonymous):

Lee shipped three packages how much did they weigh if 27 1/6 40 1/3 10 5/6 pounds

OpenStudy (texaschic101):

one weighed 27 1/6 lbs one weighed 40 1/3 lbs one weighed 10 5/6 lbs Is that correct ?

OpenStudy (anonymous):

27.16+40.3+10.83=78 1/3

OpenStudy (texaschic101):

let me check....hold on

OpenStudy (texaschic101):

27 + 40 + 10 = 77 1/6 + 1/3 + 5/6 = 1/6 + 2/6 + 5/6 = 8/6 = 1 1/3 77 + 1 1/3 = 78 1/3 you are correct :)